# Rethinking Whole-Body CT Image Interpretation: An Abnormality-Centric Approach

### v1

@@ Abstract
Automated interpretation of CT images—particularly localizing and describing abnormal findings across multi-plane and whole-body scans—remains a significant challenge in clinical radiology. This work aims to address this challenge through four key contributions: (i) On taxonomy, we collaborate with senior radiologists to propose a comprehensive hierarchical classification system, with 404 representative abnormal findings across all body regions; (ii) On data, we contribute a dataset containing over 14.5K CT images from multiple planes and all human body regions, and meticulously provide grounding annotations for over 19K abnormalities, each linked to the detailed description and cast into the taxonomy; (iii) On model development, we propose OminiAbnorm-CT, which can automatically ground and describe abnormal findings on multi-plane and whole-body CT images based on text queries, while also allowing flexible interaction through visual prompts; (iv) On benchmarks, we establish three representative evaluation tasks based on real clinical scenarios. Through extensive experiments, we show that OminiAbnorm-CT can significantly outperform existing methods on all the tasks and metrics.
@@ 1 Introduction
Computed Tomography (CT) imaging has become a cornerstone of modern medicine, with over 3 billion scans performed annually around the world. The interpretation of these scans—particularly the generation of radiology reports—plays a crucial role in guiding clinical decision-making. However, systematically identifying and characterizing abnormal findings in whole-body CT images remains a cognitively demanding and time-consuming task for radiologists rao2025multimodal .
@@
In the recent literature, artificial intelligence has started to make significant progress in CT image interpretation, particularly along two directions: first, AI models have been developed for organ-level segmentation Totalsegmentator ; jaus2023towards , that enable precise anatomical labeling and organ-grounded report generation zhang2024radgenome ; chen2025large . Yet a critical gap remains: segmenting organs alone is insufficient for clinical utility. What ultimately matters to clinicians are abnormal findings—lesions and any anomalies that inform diagnosis and treatment; second, models for automatic report generation have been trained on the recently introduced datasets with paired CT scans and corresponding radiology reports, describing clinical findings and impressions chen2024bimcv ; hamamci2024foundation . However, these datasets are typically limited to specific anatomical regions—most commonly chest CT—and often lack explicit visual grounding, which reduces explainability and increases the risk of hallucinated content. To address the above-mentioned challenges, we aim to explore the potential to develop a system that automatically detects, localizes, and describes all abnormal findings across multi-plane, whole-body CT images. This paradigm shift moves beyond regional reporting and anatomical segmentation, aligning AI interpretation more closely with the practical demands of radiological diagnosis.
Figure 1: OminiAbnorm-CT-14K is the first large-scale dataset for abnormal findings grounding and description across multi-plane and whole-body CT images. It covers 349 representative abnormal findings from 82 anatomical structures and 40 major systems or organs.
@@
One crucial barrier to developing such a system is the absence of a unified, clinically meaningful taxonomy of abnormalities, which presents a major challenge for defining task scope and establishing continuous benchmarks. In this paper, we collaborated with 7 radiologists from 3 centers (each with 10-16 years of experience) to construct a detailed hierarchical taxonomy covering 404 representative abnormal findings, that are organized across 40 major anatomical regions and 82 sub-regions.
@@
Building on this foundation, we introduce OminiAbnorm-CT-14K-the first large-scale dataset designed for abnormality grounding and description on multi-plane whole-body CT imaging. It comprises 14.5K CT images from Radiopedia radiopaedia , covering axial, coronal, and sagittal planes and diverse anatomical regions. All images and the paired reports have been rigorously reviewed by experienced radiologists. We invite 4 radiologists (with at least 7 years’ experience) to manually annotate around 19K abnormal findings on the images in the format of either bounding box or segmentation masks. All regional annotations are linked to the corresponding report descriptions, and further cast into our proposed taxonomy. These annotations cover 349 out of the 404 representative abnormal findings (86%) defined in our taxonomy across whole-body CT images. This dataset addresses a critical gap in the field, by providing high-quality, clinically relevant annotations at scale—overcoming limitations in existing datasets caused by privacy constraints and the high cost of expert annotation. All our annotations will be released to the community, while the images are accessible for research purposes after applying and paying the required fee to Radiopaedia [^1] https://radiopaedia.org/licence#data-sets.
@@
Leveraging OminiAbnorm-CT-14K, we develop OminiAbnorm-CT, a novel system for grounded CT image interpretation. By bridging a multi-modal language model with a specialized segmentation module in a bidirectional workflow, it enables dynamic invocation of the segmentation module during the text generation, and grounded abnormality interpretation based on the segmentation results. In contrast to prior approaches—which often lack grounding abilities, generalizability across planes and body regions, or flexibility in prompt handling—OminiAbnorm-CT enables automatic localization and description of abnormalities across multi-plane, whole-body CT scans. More importantly, OminiAbnorm-CT is designed to robustly interpret diverse textual references and visual prompts, enabling interactive and clinically adaptive use cases aligned with radiologists’ workflows.
For a comprehensive evaluation, we define three clinically relevant tasks: (i) grounded report generation, (ii) text-guided grounded report generation, and (iii) visual-prompted report generation. Experiment results show that OminiAbnorm-CT significantly outperforms existing baselines across all tasks, marking a substantial step toward explainable, abnormality-centric CT imaging interpretation.
@@
In summary, our contributions to advance abnormality-centric grounded CT image interpretation can be summarized as: (i) We propose a comprehensive taxonomy for abnormalities on CT images, presented in Section 2; (ii) We build OminiAbnorm-CT-14K, the first dataset for abnormality grounding and description across multi-plane whole-body CT images, detailed in Section 3; (iii) We develop a novel system OminiAbnorm-CT, for automatically localizing and reporting abnormalities, illustrated in detail in Section 4; (iv) We set up a comprehensive benchmark with three clinically representative tasks, and demonstrate the superiority of OminiAbnorm-CT in Section 5.
@@ 2 Taxonomy for Abnormal Findings
Given the diversity of abnormal findings, their complex relationships and synonyms, we first construct a comprehensive taxonomy system, to categorize all the abnormalities into representative classes.
@@
Our taxonomy is guided by four fundamental principles to ensure scientific rigor, systematic organization, and clinical utility: (i) comprehensive coverage. It should include all the clinically significant abnormalities, encompassing both lesions and the abnormal changes on anatomical structures; (ii) image-based definition. Each abnormal finding is precisely characterized by its observable features on the radiology scan, such as physical properties and morphological characteristics, that form the essential foundation for diagnosis; (iii) hierarchical organization. Findings are grouped by clinically important anatomical structures and major systems or organs, aligning with standard image interpretation protocols. While cross-regional structures (like skeletal systems) remain independent groups for conciseness; (iv) modular design. Each category represents a fundamental, independent finding that can be combined with others to express complex and multifaceted abnormalities.
@@
The taxonomy is developed through the following three stages, involving 7 radiologists with 10 to 16 years of experience from three clinical centers, (i) we establish an anatomical hierarchy framework based on standard atlases and classic textbooks for human anatomy gray2008gray ; (ii) the radiologists then systematically catalog the abnormal findings under each anatomical structure, based on their clinical experience and several authoritative textbooks adam2020grainger ; haaga2016ct ; Mandell2013CoreRadiology on medical imaging and radiological pathology; lastly, (iii) the radiologists conduct rigorous cross-validation and discussion to establish precise consensus on the definitions of controversial or ambiguous abnormalities.
We finally identified 404 representative abnormalities across 82 anatomies throughout the human body, organized within 40 major organs and systems. The details can be found in Appendix Section A.11.
3 Dataset Curation
In this section, we present OminiAbnorm-CT-14K, which establishes a data foundation for abnormality grounding and interpretation across whole-body CT images. The dataset construction pipeline is illustrated in Figure 2. Here, we start with an overview of the limitations of existing datasets in Section 3.1, which motivate the dataset construction. We then describe the data source in Section 3.2, detail the annotation procedure in Section 3.3, outline the extension to instruction-style data in Section 3.4, and conclude with a summary in Section 3.5.
3.1 Limitation in Existing Datasets
We analyze several widely used public datasets relevant to our task (summarized in Appendix Table 6) and identify the following limitations: (i) limited annotation coverage. Most datasets provide grounding annotations for only a small subset of lesion types. Despite efforts to map existing labels to our proposed taxonomy, the majority of abnormal finding categories remain unannotated; (ii) restricted imaging scope. Existing datasets are primarily limited to axial-plane CT images or focus on specific anatomical regions, overlooking the growing clinical need for interpretation across multi-plane, whole-body CT scans; (iii) insufficient image-report alignment. Many datasets lack explicit associations between localized abnormalities in the images and their corresponding descriptions in radiology reports, hindering the development of grounded, explainable systems. These limitations make existing datasets inadequate for benchmarking or developing models capable of automatic detection, localization, and description of abnormal findings in multi-plane, whole-body CT imaging.
@@ 3.2 Data Source
Following prior works wu2023towards ; zheng2024large ; wu2024mrgen , we collect raw data from Radiopaedia radiopaedia , a publicly accessible platform of peer-reviewed clinical cases contributed by clinicians worldwide, with appropriate privacy safeguards. After application, we have obtained proper permission for all data used in this study. Shown in Figure 2(a), each case includes radiology scans across various planes and anatomical regions, along with patient background information and detailed clinical reports describing findings and diagnoses. In total, we collected 46,721 CT images paired with 14,920 clinical reports.
Figure 2: Data curation overview. (a) The CT image-report pairs are collected from an open-sourced and expert-checked website; (b) and (c) Radiologists provide grounding annotation on any abnormal findings, link to their text description in reports, and categorize into the taxonomy devised by a group of senior radiologists. The annotated data are further extended to instruction data with simulated visual prompts and text queries.
@@ 3.3 Annotation Pipeline
As illustrated in Figure 2(b) and (c), our goal is to produce grounding annotations for all abnormal findings in the collected CT images, i.e., linking them to corresponding textual descriptions in clinical reports, and categorizing them using our proposed taxonomy. To reduce annotation burden, we adopt a two pre-processing strategy: (i) image-side pre-processing. Rather than annotating entire scans, we focus on the key slice and its 8 adjacent slices, as identified and verified by radiologists to be the most informative and representative; (ii) text-side pre-processing. We employ GPT-4o gpt4o to extract findings from the reports and identify abnormalities using RaTEScore zhao2024ratescore .
@@
Here, we recruited 4 clinicians with at least 7 years of radiology experience as annotators. To ensure high-quality annotations, we followed three core principles: (i) representative slices. Select the most representative slice(s) for annotation when abnormalities are present; (ii) complete annotation. Annotate all abnormal findings visible on selected slices; (iii) vision-language consistency. Only annotate abnormalities identifiable in provided slices; similarly, linked descriptions must reflect findings visible in these slices. For example, when a report mentions organ enlargement, the finding should not be annotated if enlargement cannot be confidently determined from the available slices.
@@
Throughout the annotation process, we continuously refined OminiAbnorm-CT-14K based on annotator feedback: (i) taxonomy refinement. We remove or merge 4 less representative categories, add 22 previously omitted but clinically important ones, and revise 21 categories for improved clarity and expression; (ii) long-tail mitigation. To address data imbalance, we identify the underrepresented categories and use GPT-4o to filter scans referencing these rare findings, prioritizing them for annotation. As shown in Appendix Section A.6, it has significantly mitigated the long-tail distribution.
3.4 Extension to Instruction Data
@@
As shown in Figure 2(b), we extend our annotated data to support three instruction tasks that simulate realistic clinical scenarios: (i) grounded report generation simulates a comprehensive examination. We format instructions to detect, ground, and describe all abnormal findings on CT images; (ii) text-guided grounded report generation involves responding to text queries about specific abnormalities of interest. For instance, for a patient with tuberculous empyema history, checking any abnormality related to it. We generate these queries using GPT-4o based on clinical reports, patient presentations, and medical history, simulating how radiologists approach images with prior knowledge; (iii) visual prompted report generation focuses on interpreting the marked abnormality. We create visual prompts from annotation masks using various formats, e.g., bounding box, ellipse, contour, and cropped region. This simulates a semi-automated workflow where the model elaborates the clinician-identified abnormality in details. More details are in Appendix Section A.10.

### v2

@@ Abstract
Automated interpretation of CT images—particularly localizing and describing abnormal findings across multi-plane and whole-body scans—remains a significant challenge in clinical radiology. This work aims to address this challenge through four key contributions: (i) On taxonomy, we collaborate with senior radiologists to propose a comprehensive hierarchical classification system, with 404 representative abnormal findings across all body regions; (ii) On data, we contribute a dataset containing over 14.5K CT images from multiple planes and all human body regions, and meticulously provide grounding annotations for over 19K abnormalities, each linked to the detailed description and cast into the taxonomy; (iii) On model development, we propose OmniAbnorm-CT, which can automatically ground and describe abnormal findings on multi-plane and whole-body CT images based on text queries, while also allowing flexible interaction through visual prompts; (iv) On evaluation, we establish three representative tasks based on real clinical scenarios, and introduce a clinically grounded metric to assess abnormality descriptions. Through extensive experiments, we show that OmniAbnorm-CT can significantly outperform existing methods in both internal and external validations, and across all the tasks.
Figure 1: This work introduces OmniAbnorm-CT-14K, the first large-scale dataset for grounding and describing abnormal findings in CT images. Left: OmniAbnorm-CT-14K contains 14.5K multi-plane, whole-body CT images, covering 349 representative abnormal findings across 82 anatomical structures and 40 major systems or organs. Right: Distribution of the dataset across anatomical structures and major systems or organs, with darker blue indicating higher sample density.
@@ 1 Introduction
Computed Tomography (CT) imaging has become a cornerstone of modern medicine, with over 450 million scans performed annually worldwide. The interpretation of these scans, particularly the generation of radiology reports, plays a crucial role in clinical decision-making. However, systematically identifying and characterizing abnormal findings in whole-body CT images remains a cognitively demanding and time-consuming task for radiologists [rao2025multimodal].
@@
Advances in artificial intelligence (AI) have shown promising progress in CT image interpretation, primarily along two directions: first, models developed for organ segmentation [Totalsegmentator, jaus2023towards] enable precise anatomical labeling and organ-grounded report generation [zhang2024radgenome, chen2025large]. Yet a critical gap remains: segmenting organs alone is insufficient for clinical utility. What ultimately matters to clinicians are abnormalities, lesions, and any anomalies that inform diagnosis and treatment; second, models for automatic report generation have been trained on the recent datasets with paired CT scans and radiology reports, describing findings and impressions [chen2024bimcv, hamamci2024developing]. However, these datasets are typically limited to specific anatomical regions, e.g., chest CT, and often lack explicit visual grounding, which reduces explainability and increases the risk of hallucinations. To address these challenges, we aim to develop a system that detects, localizes, and describes all abnormal findings across multi-plane, whole-body CT images automatically. This paradigm shift moves beyond regional reporting and anatomical segmentation, aligning AI interpretation more closely with practical radiological demands.
@@
One crucial barrier to developing such a system is the absence of a unified, clinically meaningful taxonomy of abnormalities, which hinders the definition of task scope and continuous benchmarks. In this paper, we collaborated with 7 senior radiologists from 3 centers (each with 10-16 years of experience) to construct a hierarchical taxonomy covering 404 representative abnormal findings, organized across 40 major anatomical regions and 82 sub-regions.
@@
Based on the taxonomy, we introduce OmniAbnorm-CT-14K, the first large-scale dataset designed for abnormality grounding and description on multi-plane whole-body CT imaging. It comprises 14.5K CT images from Radiopedia [radiopaedia], covering axial, coronal, and sagittal planes and diverse anatomical regions. All images and the paired reports are contributed and rigorously reviewed by worldwide radiologists. We invite 4 board-certified radiologists to manually annotate around 19K abnormal findings on the images in the format of either bounding box or segmentation masks. All regional annotations are linked to the corresponding report descriptions, and further cast into our proposed taxonomy. These annotations cover 349 out of the 404 representative abnormal findings (86%) defined in our taxonomy across whole-body CT images. By providing high-quality, clinically relevant annotations at scale, this dataset addresses the critical data limitations in the field, caused by privacy constraints or high annotation cost. All our annotations will be released to the community, while the images can be accessed for research purposes after application [^1] https://radiopaedia.org/licence#data-sets.
@@
Leveraging this dataset, we develop OmniAbnorm-CT, a novel system for grounded CT image interpretation. By bridging a vision language model (VLM) with the specialized segmentation module in a bidirectional workflow, it enables the VLM to dynamiclly invoke the segmentation module during generation, and generate abnormality interpretation grounded on the segmentation results. In contrast to prior approaches, which often lack grounding abilities, generalizability across planes and body regions, or flexibility in prompt handling, OmniAbnorm-CT enables automatic localization and description of abnormalities across multi-plane, whole-body CT images. Moreover, OmniAbnorm-CT is designed to robustly interpret diverse textual and visual prompts, enabling interactive and adaptive use cases aligned with radiologists’ clinical workflows.
For a comprehensive evaluation, we first define three clinically relevant tasks: (i) grounded report generation to detect, localize, and describe all findings; (ii) text-guided grounded report generation to detect, localize, and describe abnormalities based on textual queries about specific clinical concerns, and (iii) visual-prompted report generation to describe a particular abnormality indicated by a visual prompt. Then, we propose a clinically grounded metric for report evaluation, named AbnormRubric. In contrast to existing metrics relying on simple matching [bleu, rouge, banerjee2005meteor] or overall semantic similarity [zhangbertscore, zhao2024ratescore, delbrouck2024radgraph], AbnormRubric customizes the rubric for each abnormality category, ensure alignment with clinical judgement. Implemented using a large language model (LLM) as a judge, it explicitly quantifies recall and precision in abnormality detection, along with the descriptive accuracy of abnormalitis across four key radiological attributions: location, morphology, density, and size. Experiment results show that OmniAbnorm-CT significantly outperforms existing baselines across all tasks, marking a substantial step toward explainable, abnormality-centric CT imaging interpretation.
@@
In summary, our contributions to advance abnormality-centric grounded CT image interpretation are: (i) A comprehensive taxonomy for abnormalities on CT images, presented in Section 3; (ii) OmniAbnorm-CT-14K, the first dataset for abnormality grounding and description across multi-plane whole-body CT images, detailed in Section 4; (iii) OmniAbnorm-CT, a system for automatically localizing and reporting abnormalities, illustrated in Section 5; (iv) A benchmark with three representative tasks and a clinically grounded metrics for report evaluation, demonstrating the superiority of OmniAbnorm-CT in Section 6.
2 Related Work
Medical image segmentation aims to delineate clinically meaningful regions on medical images. For radiology, most datasets focus on annotations for important organs and anatomical structures [Totalsegmentator, abdomenatlas, FLARE, AMOS22] or lesions [heller2021kidney, MSD, BraTS2023GLI, de2024uls23, yan2018deep]. The past decade witnessed both remarkable success in specialist segmentation models [nnUNET, UNet, zhou2023nnformer, zhou2019unet++] and a recent paradigm shift towards building generalist models [MedSAM, zhao2024foundation, zhao2025large, du2024segvol]. However, few works contribute to the grounding of abnormal image findings, a broader spectrum including abnormal organs, potential lesions, and more generally, any anomalies meaningful for clinical decision-making. A concurrent work [baharoon2025rexgroundingct] provide segmentation for 14 types of abnormalities, while limited to chest CT.
Medical image report generation aims to faithfully interpret the medical images with formal text, which is indispensable for numerous clinic procedures [rao2025multimodal]. Existing datasets with image-report pairs are limited to specific modalities and regions, such as chest X-ray [johnson2019mimic], brain MRI [lei2024autorg] and chest CT [hamamci2024developing, chen2024bimcv], overlooking the demands for multi-plane whole-body CT report generation. Recent progress [zhang2024radgenome, chen2025large, lei2024autorg, chen2024vision] on grounded report generation is limited to organ-level segmentation and description, failing to localize and interpret fine-grained abnormality.
Large visual-language models (LVLM) for medicine are built on large-scale multi-modal medical data, demonstrating exceptional performance and generalization capabilities across diverse tasks [wu2023towards, moor2023med, zhang2024generalist, li2023llava, meddr]. Despite their potential application to multi-plane whole-body CT images, most existing approaches fail to provide grounded evidence during generation simultaneously, leading to poor explainability and hallucinations that are more difficult to detect.
@@ 3 Taxonomy for Abnormal Findings
Given the diversity, complex relationship, and synonyms of abnormalities, we first construct a taxonomy system to categorize all the abnormalities into representative classes.
@@
Our taxonomy is guided by four fundamental principles to ensure scientific rigor, systematic organization, and clinical utility: (i) comprehensive coverage. It should include all the clinically significant abnormalities, encompassing both lesions and the abnormal changes on anatomical structures; (ii) image-based definition. Each abnormal finding is precisely characterized by its observable features on the radiology image, such as physical properties and morphological characteristics that form the essential foundation for diagnosis; (iii) hierarchical organization. Findings are grouped by anatomical structures and major systems or organs, aligning with standard image interpretation protocols. While cross-regional structures (like skeletal systems) remain independent groups for conciseness; (iv) modular design. Each category represents a fundamental, independent finding that can be combined with others to express complex and multifaceted abnormalities.
@@
The taxonomy is developed by 7 senior radiologists with at least 10-year experience from 3 centers, following three stages: (i) an anatomical hierarchy framework is establish based on standard atlases and authoritative human anatomy textbooks [gray2008gray]; (ii) abnormal findings are systematically cataloged under each anatomical structure, based on the clinical experience and authoritative textbooks [adam2020grainger, haaga2016ct, Mandell2013CoreRadiology] on radiological imaging and pathology; (iii) the radiologists conduct rigorous cross-validation and discussion to establish precise consensus on the definitions of controversial or ambiguous abnormalities. We finally identified 404 representative abnormalities across 82 anatomies throughout the human body, organized within 40 major organs or systems.
4 OmniAbnorm-CT-14K
Figure 2: Data curation overview. (a) The image-report pairs are collected from an open-sourced and expert-checked website; (b) and (c) Radiologists provide grounding annotation on any abnormalities, link to their text description in reports, and categorize into the taxonomy devised by senior radiologists. The annotation is further extended to instruction data with simulated visual prompts and text queries.
@@ 4.1 Data Source
Following prior works [wu2023towards, zheng2024large, wu2024mrgen], we collect raw data from Radiopaedia [radiopaedia], a publicly accessible platform of peer-reviewed clinical cases contributed by clinicians worldwide, with appropriate privacy safeguards. After application, we have obtained proper permission for all data used in this study. Shown in Figure 2(a), each case includes radiology images across various planes and anatomical regions, along with patient background information and clinical reports describing findings and diagnoses. In total, we collected 46,721 CT images paired with 14,920 clinical reports.
@@ 4.2 Annotation Pipeline
As illustrated in Figure 2(b) and (c), our goal is to produce grounding annotations for all abnormal findings in the collected CT images, i.e., linking them to corresponding textual descriptions in clinical reports, and categorizing them using our proposed taxonomy. To reduce annotation burden, we adopt a two pre-processing strategy: (i) image-side pre-processing. Rather than annotating entire scans, we focus on one key slice and its 8 adjacent slices, selected by original uploading radiologists to be the most informative and representative; (ii) text-side pre-processing. We employ GPT-4o [gpt4o] to extract findings from the reports and identify abnormalities using RaTEScore [zhao2024ratescore].
@@
Then, we recruited 4 board-certified radiologists as annotators through a selection phase: candidates each annotated 100 test cases; senior radiologists then selected the top four for full-scale annotation. To ensure high-quality annotations, we followed three core principles: (i) representative slices. Select the most representative slice(s) for annotation when abnormalities are present; (ii) complete annotation. Annotate all abnormal findings visible on selected slices; (iii) vision-language consistency. Only annotate abnormalities identifiable in provided slices; similarly, linked descriptions must reflect findings visible in these slices. For example, when a report mentions organ enlargement, the finding should not be annotated if enlargement cannot be confidently determined from the available slices.
@@
Throughout the annotation, senior radiologists provided real-time consultation and guidance to resolve uncertain cases from the annotators. Meanwhile, we continuously refined OmniAbnorm-CT-14K based on annotator feedback: (i) taxonomy refinement. We remove or merge 4 less representative categories, add 22 previously omitted but clinically important ones, and revise 21 categories for improved clarity and expression; (ii) long-tail mitigation. To address data imbalance, we identify the underrepresented categories and use GPT-4o [gpt4o] to filter scans referencing these rare findings, prioritizing them for annotation. More details about the long-tail mitigation can be found in the Appendix.
4.3 Instruction Data Construction
@@
As shown in Figure 2(b), we extend our annotated data to support three instruction tasks that simulate realistic clinical scenarios: (i) grounded report generation simulates a comprehensive examination. We format instructions to detect, ground, and describe all abnormal findings on CT images; (ii) text-guided grounded report generation involves responding to text queries about specific abnormalities of interest. For instance, for a patient with tuberculous empyema history, checking any abnormality related to it. We generate these queries using GPT-4o based on clinical reports, patient presentations, and medical history, simulating how radiologists approach images with prior knowledge; (iii) visual prompted report generation focuses on interpreting the marked abnormality. We create visual prompts from annotation masks using various formats, e.g., bounding box, ellipse, contour, and cropped region. This simulates a semi-automated workflow where the model elaborates on the clinician-identified abnormality in detail. More details about the instruction data are in the Appendix.